OpenStudy (anonymous):

WILL AWARD MEDAL! A student is assessing the correlation between the average number of hours of playing video games in a week and average score on a math test for the students of a class. The table below shows the data: Number of hours of playing video games (x) 0 0.5 1 1.5 2 2.5 3 3.5 4 4.5 Score on math test (y) 90 88 86 84 82 80 78 76 74 72 Part A: Is there any correlation between the average number of hours of playing video games in a week and average score on the math test for the students of a class? Justify your answer.

OpenStudy (anonymous):

Part B: Write a function which best fits the data. Part C: What does the slope and y-intercept of the plot indicate?

OpenStudy (anonymous):

@Alexandra675

OpenStudy (anonymous):

@Clalgee

OpenStudy (anonymous):

all i know is that there is a pattern between the numbers

OpenStudy (anonymous):

the more hours on a video game the worse the score

OpenStudy (anonymous):

so there is a correlation every half an hour there score decreases by 2

OpenStudy (anonymous):

do ya got it

OpenStudy (anonymous):

@cassious

OpenStudy (anonymous):

okay so that would be part A. what are your thoughts for Part B and C?

OpenStudy (anonymous):

im not so sure about those

OpenStudy (anonymous):

:/ well thanks anyway

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@CrazyMexican816 help @cassious

OpenStudy (anonymous):

@marissalovescats

OpenStudy (marissalovescats):

I'd say yes there is a correlation because as you can see the longer people play the games, the lower their test scores go

OpenStudy (anonymous):

i already said that he needs b and c

OpenStudy (marissalovescats):

Oh oops sorry

OpenStudy (anonymous):

lol

OpenStudy (marissalovescats):

Well for part A specifically you should say that each time they play the game for 0.5 hours more, their test score goes down 2 points.

OpenStudy (marissalovescats):

For Part B you need what I just wrote for Part A. I'd say y-2=x+0.5

OpenStudy (marissalovescats):

For Part C the slope shows the number of hours they are going up and the y intercept shows how their test scores are changing when they play more

OpenStudy (anonymous):

THANK YOU!

OpenStudy (amistre64):

thats a little off ....

OpenStudy (marissalovescats):

Or I'm wrong :P

OpenStudy (marissalovescats):

Happens to everyone some time around

OpenStudy (amistre64):

since both x and y have a constant rate of change, they can be modeled with the equation of a line, but your parts of off.

OpenStudy (amistre64):

the slope of the line is: change in y --------- change in x and is conventionally named by the variable m once we know the slope, we can form the function using point slope format as you attemted: take a known point (a,b) to construct the equation as: y - b = m(x-a), or more familiarly as: y = mx +(-ma+b)

OpenStudy (amistre64):

y-2=x+0.5 has a slope of 1, and is attached to the point (.5,2) which does not correspond with the data given.

OpenStudy (amistre64):

*the point (-.5,2) at best

OpenStudy (anonymous):

so that would be for part B,correct?

OpenStudy (amistre64):

that is the basic process for part B yes part A is asking you to define the slope part B is asking you to define the line equation part C is asking you to use the line equation for find the value when x=0